# Image Matching Filtering and Refinement by Planes and Beyond

### v3

@@ 2.3.1 Standalone pipeline modules
Deep learning evolution has provided a boost to image matching as well as other computer vision research fields thanks to the evolution of the deep architectures, hardware and software computational capability and design, and the huge amount of datasets now available for the training. Hybrid pipelines are raised first, replacing handcrafted descriptors with deep ones, as the feature descriptor module is the most natural and immediate to be replaced in the pipeline evolution. SOTA in this sense is represented by the Hard Network (HardNet) [80] and similar networks [81, 82] trained on contrastive learning with hard mining. For what concerns the keypoint extraction, better results with respect to the handcrafted counterparts were achieved by driving the network to mimic handcrafted detector design as for Key.Net [52]. Deep learning has been successfully applied to design standalone patch normalization modules to estimate the reference patch orientation [83, 84] or more complete affine transformations as in the Affine Network [84]. The mentioned deep modules have been employed successfully in many hybrid matching pipelines, especially in conjunction with SIFT [11] or Harris corners [85].
@@ 2.4 Joint detector & descriptor networks
A relevant characteristic in deep image matching which made it able to exceed handcrafted image matching pipelines was gained by allowing concurrently optimizing both the detector & descriptor. SuperPoint [53] integrates both of them in a single network and does not require a separate yet consecutive training of these as [86]. A further design feature that enabled SuperPoint to become a reference in the field is the use of homographic adaptation, which consists of the use of planar homographies to generate corresponding images during the training, to allow self-supervised training. Further solutions were proposed by architectures similar to SuperPoint as in [87]. Moreover, the Accurate and Lightweight Keypoint Detection and Descriptor Extraction (ALIKE) [88] implements a differentiable Non-Maximum Suppression (NMS) for the keypoint selection on the score map, and the DIScrete Keypoints (DISK) [89] employs reinforcement learning. ALIKE design is further improved by A LIghter Keypoint and descriptor Extraction network with Deformable transformation (ALIKED) [13] using deformable local features for each sparse keypoint instead of dense descriptor maps, resulting in saving computational budget. Alternative rearrangements in the pipeline structure to use the same sub-network for both detector & descriptor (detect-and-describe) [90], or to extract descriptors and then keypoints (describe-then-detect) [91] instead of the common approach (detect-then-describe) have been also investigated. More recently, the idea of decoupling the two blocks by formulating the matching problems in terms of 3D tracks in large-scale SfM has been applied in Detect, Don’t Describe–Describe, Don’t Detect (DeDoDe) [92] and DeDoDe v2 [93] with interesting results.
@@ 2.4.1 Full end-to-end architectures
Full end-to-end deep image matching architectures arose with SuperGlue [94], which incorporates a proper matching network exploiting self and cross attentions through transformers to jointly encode keypoint positions and visual data into a graph structure, and a differentiable form of the Sinkhorn algorithm to provide the final assignments. SuperGlue can represent a breakthrough in the areas and has been extended by the recent LightGlue [12] improving its efficiency, accuracy and training process. Detector-free end-to-end image matching architectures have been also proposed, which avoid explicitly computing a sparse keypoint map, considering instead a semi-dense strategy. The Local Feature TRansformer (LoFTR) [14] is a competitive realization of these matching architectures, which deploys a coarse-to-fine strategy using self and cross attention layers on dense feature maps and has been employed as a base for further design extensions [15, 16]. Notice that end-to-end image matching networks are not fully rotational invariant and can handle only relatively small scene rotations, except specific architectures designed with this purpose [95, 96, 97]. Coarse-to-fine strategy following LoFTR design has been also applied to hybrid canonical pipelines with only deep patch and descriptor modules, leading to matching results comparable in challenging scenarios at an increased computation cost [50]. More recently, dense image matching networks that directly estimate the scene as dense point maps [17] or employing Gaussian process [18], achieve SOTA results in many benchmarks but are more computationally expensive than sparse or semi-dense methods.
@@ 2.4.2 Outlier rejection with geometric constraints
The first effective deep matching filter module was achieved with the introduction of the context normalization [34], which is the normalization by mean and standard deviation, that guarantees to preserve permutation equivalence. The Attentive Context Network (ACNe) [35] further includes local and global attention to exclude outliers from context normalization, while the Order-Aware Network (OANet) [36] adds additional layers to learn how to cluster unordered sets of correspondences to incorporate the data context and the spatial correlation. The Consensus Learning Network (CLNet) [37] prunes matches by filtering data according a local-to-global dynamic neighborhood consensus graphs in consecutive rounds. The Neighbor Consistency Mining Network (NCMNet) [38] improves CLNet architecture by considering different neighborhoods in the feature space, the keypoint coordinate space, and a further global space that considers both previous ones altogether with cross-attention. More recently, the Multiple Sparse Semantics Dynamic Graph Network (MS2DG-Net) [39] designs the filtering in terms of a neighborhood graph through transformers [98] to capture semantics similar structures in the local topology among correspondences. Unlike previous deep approaches, ConvMatch[40] builds a smooth motion field by making use of Convolutional Neural Network (CNN) layers to verify the consistency of the matches. DeMatch [41] refines ConvMatch motion field estimation for a better accommodation of scene disparities by decomposing the rough motion field into several sub-fields. Finally, notice that deep differentiable RANSAC modules have been investigated as well [99, 100].
@@ 2.4.3 Keypoint refinement
The coarse-to-fine strategy trend has pushed the research of solutions focusing on the refinement of keypoint localization as a standalone module after the coarse assignment of the matches. Worth mentioning is Pixel-Perfect SfM [101] that learns to refine multi-view tracks on their photometric appearance and the SfM structure, while Patch2Pix [42] refines and filters patches exploiting image dense features extracted by the network backbone. Similar to Patch2Pix, Keypt2Subpx [43] is a recent lightweight module to refine only matches requiring as input the descriptors of the local heat map of the related correspondences beside keypoints position. Finally, the Filtering and Calibrating Graph Neural Network (FC-GNN) [44] is an attention-based Graph Neural Network (GNN) jointly leveraging contextual and local information to both filter and refine matches. Unlike Patch2Pix and Keypt2Subpx, does not require feature maps or descriptor confidences as input, as also the proposed MOP+MiHo+NCC. Similarly to FC-GNN, the proposed approach uses both global context and local information to refine matches, respectively to warp and check photometric consistency, but differently from FC-GNN it currently only employs global contextual data since does not implement any feedback system to discard matches on patch similarity after the alignment refinement.
@@ 3.1.1 Preliminaries
MOP assumes that motion flow within the images can be piecewise approximated by local planar homography.
Representing with an abuse of notation the reprojection of a point through a planar homography with non-singular as , a match within two keypoints and is considered an inlier based on the maximum reprojection error
@@
A proper scene planar transformation must be quasi-affine to preserve the convex hull [45] with respect to the minimum sampled model generating in RANSAC so the inlier set effectively employed is
where is the set of matches in satisfying the quasi-affininity property with respect to the RANSAC minimum model of described later in Eq. 12.
@@ 3.1.2 Main loop
Starting with , and the initial set of input matches provided by the image matching pipeline, MOP at each iteration looks for the best planar homography compatible with the current match set according to a relaxed inlier threshold px. When the number of inliers provided by is less than the minimum required amount, set to , i.e.
the current homography is discarded and the failure counter is incremented without updating the iteration counter , the next match set and the output list . Otherwise, MOP removes the inlier matches of for a threshold for the next iteration
@@ 3.1.3 Inside RANSAC
Furthermore, being the corresponding homography derived by the normalized 8-point algorithm [45], the smaller singular value found by solving through SVD the homogeneous system must be relatively greater than zero for a stable solution not close to degeneracy. According to these observations, the minimum singular value is constrained to be greater than 0.05. Note that an absolute threshold is feasible since data are normalized before SVD.
Next, matches in must satisfy quasi-affinity with respect to the derived homography . This can be verified by checking that the sign of the last coordinate of the reprojection in non-normalized homogenous coordinates are the same for , i.e. for all the four keypoint in the first image. This can be expressed as
@@
Lastly, to speed up the RANSAC search, a buffer with is globally maintained in order to contain the top discarded sub-optimal homographies encountered within a RANSAC run. For the current match set if is the current best RANSAC model maximizing the number of inliers, are ordered by the number of inliers excluding those compatible with the previous homographies in the buffer, i.e. it must holds that
@@ 3.1.4 Homography assignment
Next, a homography must be associated with each survived match , which will be used for patch normalization.
A possible choice is to assign the homography which gives the minimum reprojection error, i.e.
However, this choice is likely to select homographies corresponding to narrow planes in the image with small consensus sets, which leads to a low reprojection error but more unstable and prone to noise assignment.
@@
which points to more stable and robust homographies but can lead to incorrect patches when the corresponding reprojection error is not among the minimum ones.
The homography actually chosen for the match association provides a solution within the above. Specifically, for a match , defining as the median of the number of inliers of the top 5 compatible homographies according to the number of inliers for , is the compatible homography with the minimum reprojection error among those with an inlier number at least equal to
Figure 1 shows an example of the achieved solution directly in combination with MiHo homography estimation described hereafter in Sec. 3.2, since MOP or MOP+MiHo in visual cluster representation does not present appreciable differences. Keypoint belonging to discarded matches are marked with black diamonds, while the clusters highlighted by other combinations of markers and colors indicate the resulting filtered matches with the selected planar homography . Notice that clusters are not designed to highlight scene planes but points that move according to the same homography transformation within the image pair.
@@ 3.2.1 Rationale
The idea behind MiHo is to break the homography in the middle by two homographies and such that so that the patch on gets deformed by less than , and likewise the patch on gets less deformed by than . This means that visually on the reprojected patch a unit area square should remain almost similar to the original one. As this must hold from both the images, the deformation error must be distributed almost equally between the two homographies and . Moreover, since interpolation degrades with up-sampling, MiHo aims to balance the down-sampling of the patch at finer resolution and the up-sampling of the patch at the coarser resolution.
@@ 3.2.2 Implementation
While a strict analytical formulation of the above problem which leads to a practical satisfying solution is not easy to derive, the heuristic found in [51] can be exploited to modify RANSAC within MOP to account for the required constraints. Specifically, each match is replaced by two corresponding matches and where is the midpoint within the two keypoints
@@
leading to two concurrent homographies and to be verified simultaneously with the inlier set in analogous way to Eq. 3
@@
All the other MOP steps defined in Sec. 3.1, including RANSAC degeneracy checks, threshold adaptation, and final homography assignment, follow straightly in an analogous way. Overall, the principal difference is that a pair of homography is obtained instead of the single homography to be operated simultaneously on split match sets and instead of the original one .
Figure 2 shows the differences in applying MiHo to align to planar scenes with respect to directly reproject in any of the two images considered as reference. It can be noted that the distortion differences are overall reduced, as highlighted by the corresponding grid deformation. Moreover, MiHo homography strongly resembles an affine transformation or even a similarity, i.e. a rotation and scale change only, imposing in some sense a tight constraint on the transformation than the base homography. To balance this additional restrain with respect to the original MOP, in MOP+MiHo the minimum requirement amount of inliers for increasing the failure counter defined by Eq. 4 has been relaxed after experimental validation from to .
@@ 3.2.3 Translation invariance
where is a fixed translation vector is added to all the corresponding original midpoints . Overall, translation keypoints in the respective images have the effect of adding a fixed translation to the corresponding midpoint. Hence MiHo is invariant to translations since the normalization employed by the normalized 8-point algorithm is invariant to similarity transformations [45], including translations.

### v4

@@ 2.3.1 Standalone pipeline modules
Deep learning evolution has provided a boost to image matching as well as other computer vision research fields thanks to the evolution of the deep architectures, hardware and software computational capability and design, and the huge amount of datasets now available for the training. Hybrid pipelines raised first, replacing handcrafted descriptors with deep ones, as the feature descriptor module is the most natural and immediate to be replaced in the pipeline evolution. SOTA in this sense is represented by the Hard Network (HardNet) [hardnet] and similar networks [sosnet, hynet] trained on contrastive learning with hard mining. For what concerns the keypoint extraction, better results with respect to the handcrafted counterparts were achieved by driving the network to mimic handcrafted detector design as for Key.Net [keynet]. Deep learning has been successfully applied to design standalone patch normalization modules to estimate the reference patch orientation [learning_orientation, affnet] or more complete affine transformations as in the Affine Network [affnet]. The mentioned deep modules have been employed successfully in many hybrid matching pipelines, especially in conjunction with SIFT [imw2020] or Harris corners [hz_pipeline].
@@ 2.4 Joint detector & descriptor networks
A relevant characteristic in deep image matching which made it able to exceed handcrafted image matching pipelines was gained by allowing concurrently optimizing both the detector & descriptor. SuperPoint [superpoint] integrates both of them in a single network and does not require a separate yet consecutive training of these as [lift]. A further design feature that enabled SuperPoint to become a reference in the field is the use of homographic adaptation, which consists of the use of planar homographies to generate corresponding images during the training, to allow self-supervised training. Further solutions were proposed by architectures similar to SuperPoint as in [r2d2]. Moreover, the Accurate and Lightweight Keypoint Detection and Descriptor Extraction (ALIKE) [alike] implements a differentiable Non-Maximum Suppression (NMS) for the keypoint selection on the score map, and the DIScrete Keypoints (DISK) [disk] employs reinforcement learning. ALIKE design is further improved by A LIghter Keypoint and descriptor Extraction network with Deformable transformation (ALIKED) [aliked], which exploits deformable local features for each sparse keypoint instead of dense descriptor maps, saving computational budget. Alternative rearrangements in the pipeline structure to use the same sub-network for both detector & descriptor (detect-and-describe) [d2net], or to extract descriptors and then keypoints (describe-then-detect) [d2d] instead of the common approach (detect-then-describe), have been also investigated. More recently, the idea of decoupling the two blocks by formulating the matching problems in terms of 3D tracks in large-scale SfM has been applied in Detect, Don’t Describe–Describe, Don’t Detect (DeDoDe) [dedode] and DeDoDe v2 [dedodev2] with interesting results.
@@ 2.4.1 Full end-to-end architectures
Full end-to-end deep image matching architectures arose with SuperGlue [superglue], which incorporates a proper matching network exploiting self and cross attentions through transformers to jointly encode keypoint positions and visual data into a graph structure, and a differentiable form of the Sinkhorn algorithm to provide the final assignments. SuperGlue can represent a breakthrough in the areas and has been extended by the recent LightGlue [lightglue] improving efficiency, accuracy and the training process. Detector-free end-to-end image matching architectures have been also proposed, which avoid explicitly computing a sparse keypoint map, considering instead a semi-dense strategy. The Local Feature TRansformer (LoFTR) [loftr] is a competitive realization of these matching architectures, which deploys a coarse-to-fine strategy using self and cross attention layers on dense feature maps, and has been employed as a base for further design extensions [quadtree, matchformer]. Notice that end-to-end image matching networks are not fully rotational invariant and can handle only relatively small scene rotations, except specific architectures designed with this purpose [se2loftr, steerers, strek]. Coarse-to-fine strategy following LoFTR design has been also applied to hybrid canonical pipelines with only deep patch and descriptor modules, leading to matching results comparable in challenging scenarios at an increased computation cost [slime]. More recently, dense image matching networks that directly estimate the scene as dense point maps, as Matching And Stereo 3D Reconstruction (MASt3R) [mast3r] and others [dust3r, vggt], or employing Gaussian process as the Robust Matching (RoMa) [roma], achieve SOTA results in many benchmarks but are more computationally expensive than sparse or semi-dense methods.
@@ 2.4.2 Outlier rejection with geometric constraints
The first effective deep matching filter module was achieved with the introduction of the context normalization [learning_correspondences], which is the normalization by mean and standard deviation, that guarantees to preserve permutation equivalence. The Attentive Context Network (ACNe) [acne] further includes local and global attention to exclude outliers from context normalization. Next, the Order-Aware Network (OANet) [oanet] adds additional layers to learn how to cluster unordered sets of correspondences to incorporate the data context and the spatial correlation. The Consensus Learning Network (CLNet) [clnet] prunes matches by filtering data according a local-to-global dynamic neighborhood consensus graphs in consecutive rounds. The Neighbor Consistency Mining Network (NCMNet) [ncmnet] improves the CLNet architecture by considering different neighborhoods in the feature space, the keypoint coordinate space, and a further global space that considers both previous ones altogether with cross-attention. More recently, the Multiple Sparse Semantics Dynamic Graph Network (MS2DG-Net) [ms2dgnet] designs the filtering in terms of a neighborhood graph through transformers [transformer] to capture semantics similar structures in the local topology among correspondences. Unlike previous deep approaches, ConvMatch[convmatch] builds a smooth motion field by making use of Convolutional Neural Network (CNN) layers to verify the consistency of the matches. DeMatch [dematch] refines ConvMatch motion field estimation for a better accommodation of scene disparities by decomposing the rough motion field into several sub-fields. Finally, notice that deep differentiable RANSAC modules have been investigated as well [ngransac, generalizedransac].
@@ 2.4.3 Keypoint refinement
The coarse-to-fine strategy trend has pushed the research of solutions focusing on the refinement of keypoint localization as a standalone module after the coarse assignment of the matches. Worth mentioning is Pixel-Perfect SfM [pixelperfectsfm] that learns to refine multi-view tracks on their photometric appearance and the SfM structure, while Patch2Pix [patch2pix] refines and filters patches exploiting image dense features extracted by the network backbone. Similar to Patch2Pix, Keypt2Subpx [keypt2subpx] is a recent lightweight module to refine matches only which requires as input, beside keypoint positions, the descriptors of the local heat map of the related correspondences.
Another approach related to ASIFT was proposed in [monoref], where the planes on the scene are extracted by a single image deep-estimation network independently for each image, so as to rectified the corresponding image regions accordingly before extracting keypoints and patches. Unlike this approach, the proposed MOP-based approaches do not require real planes in the scene and handle patch planar deformations. Moreover, MOP-based approaches do not rely on and need a single image deep-estimation network in order to extract the scene 3D structure. That network can be unable to resolve ambiguities in common planar scenes leading to failure cases not present in base pipelines as well as in our approach. Unlike [monoref] which finds planes and then extract keypoints, the proposed approach finds planes from the provided keypoints and then refines these latter ones.
Finally, the Filtering and Calibrating Graph Neural Network (FC-GNN) [fcgnn] is an attention-based Graph Neural Network (GNN) jointly leveraging contextual and local information to both filter and refine matches. Unlike Patch2Pix and Keypt2Subpx, FC-GNN does not require feature maps or descriptor confidences as input, as also the proposed MOP-based approach. On one hand, similarly to FC-GNN, the proposed approach uses both global context and local information to refine matches, respectively to warp and check photometric consistency. On the other hand, differently from FC-GNN, the proposed approach currently only employs global contextual data since it does not implement any feedback system to discard matches on patch similarity after the alignment refinement.
@@ 3.1.1 Preliminaries
MOP assumes that motion flows within the images can be piecewise approximated by local planar homography.
Representing with an abuse of notation the reprojection of a point through a planar homography as , where is non-singular, a match within two keypoints and is considered an inlier based on the maximum reprojection error
@@
A proper scene planar transformation must be quasi-affine to preserve the convex hull [multiview] with respect to the minimum sampled model generating in RANSAC. The the inlier set effectively employed is
where is the set of matches in satisfying the quasi-affinity property with respect to the RANSAC minimum model of , described later in Eq. 12.
@@ 3.1.2 Main loop
Starting with , and the initial set of input matches provided by the image matching pipeline, at each iteration MOP looks for the best planar homography compatible with the current match set according to a relaxed inlier threshold px. When the number of inliers provided by is less than the minimum required amount, set to , i.e.
the current homography is discarded and the failure counter is incremented without updating the iteration counter , the next match set and the output list . Otherwise, MOP removes the inlier matches of for a given threshold in the next iteration
@@ 3.1.3 Inside RANSAC
Furthermore, being the corresponding homography derived by the normalized Direct Linear Transform (DLT) algorithm [multiview], the smaller singular value found by solving through SVD the homogeneous system must be relatively greater than zero for a stable solution not close to degeneracy. According to these observations, the minimum singular value is constrained to be greater than 0.05. Note that an absolute threshold is feasible since data are normalized before SVD.
Next, matches in must satisfy quasi-affinity with respect to the derived homography . This can be verified by checking that the sign of the last coordinate of the reprojection in non-normalized homogenous coordinates is the same for , i.e. for all the four keypoint in the first image. This constraint can be expressed as
@@
Lastly, to speed up the RANSAC search, a buffer with is globally maintained in order to contain the top discarded sub-optimal homographies encountered within a RANSAC run. For the current match set , if is the current best RANSAC model maximizing the number of inliers, are ordered by the number of inliers excluding those compatible with the previous homographies in the buffer, i.e. it must holds that
@@ 3.1.4 Homography assignment
The final set of pruned matches can be directly employed in the next step of the pipeline when no keypoint refinement is required, accomplishing the MOP filtering task.
Otherwise, a homography must be associated with each survived match , which will be used for patch normalization in order to proceeds with NCC template matching. A possible choice is to assign the homography which gives the minimum reprojection error, i.e.
However, this choice is likely to select homographies corresponding to narrow planes in the image with small consensus sets, which leads to a low reprojection error but also to a more unstable and prone to noise assignment.
@@
which points to more stable and robust homographies, but can lead to incorrect patches when the corresponding reprojection error is not among the lowest ones.
The homography actually chosen for the match association provides an intermediate solution between the previous ones. Specifically, for a match , defining as the median of the number of inliers of the top 5 compatible homographies according to the number of inliers for , is defined as the compatible homography with the minimum reprojection error among those with an inlier number at least equal to
Figure 1 shows an example of the achieved solution directly in combination with MiHo homography estimation described hereafter in Sec. 3.2. MOP or MOP+MiHo cluster representation does not present visually appreciable differences. Keypoint belonging to discarded matches are marked with black diamonds, while the clusters highlighted by other combinations of markers and colors indicate the resulting filtered matches with the selected planar homography . Notice that clusters are not designed to highlight scene planes but points that move according to the same homography transformation within the image pair.
@@ 3.2.1 Rationale
MiHo is aimed to alleviate the above issue by providing symmetric patch rectification for matches and error threshold consistency. The idea behind MiHo is to break the homography in the middle by two homographies and such that . The assumption is that the patch on gets deformed by less than by , and likewise the patch on gets deformed by less than by . This means that visually on the reprojected patch a unit area square should remain almost similar to the original one. As this must hold from both the images, the deformation error must be distributed almost equally between the two homographies and . Moreover, since interpolation degrades with up-sampling, MiHo aims to balance the down-sampling of the patch at finer resolution and the up-sampling of the patch at the coarser resolution.
@@ 3.2.2 Implementation
While a strict analytical formulation of the above problem which should lead to a practical satisfying solution is not easy to derive, the heuristic found in [average_homography] can be exploited to modify the base RANSAC employed by MOP described in Sec. 3.1.3 to account for the required constraints. Specifically, each match is replaced by two corresponding matches and where is the midpoint within the two keypoints
@@
leading to two concurrent homographies and to be verified simultaneously with the inlier set in analogous way to Eq. 3, i.e.
@@
All the other MOP steps detailed in Sec. 3.1.3, including RANSAC degeneracy checks, threshold adaptation, and final homography assignment, follow straightly in an analogous way.
Overall, the principal difference to the base MOP is that two concurrent RANSAC steps are executed, corresponding to the replacement of the actual match set by the two linked match sets and . Analogously, the sampling minimal set is replaced by and . This leads to obtain pairs of homography instead of the single homography inside the RANSAC loop, so that a match of is to be considered inlier if the corresponding pair of matches are respectively inliers for and .
3.2.3 Middle and half homographies
MiHo shares some similarities with the half homography concept described in [VSAC] with the purpose of adjust the matched keypoints after the final homography has been estimated. Specifically, the decomposition of the original homography matrix into the half homography matrix provided by the square root of is employed as replacement of both and . The VSAC half homography can be used to directly split the homographies computed by the base MOP into pairs of homographies instead of MiHo.
Figure 2-3 show the differences in applying MiHo to align to planar scenes with respect to directly reproject in any of the two images considered as reference, or using VSAC half homography. More examples can be found in the supplemental material2. With respect to the direct reprojections, it can be noted that MiHo distortions are overall reduced, as highlighted by the corresponding grid deformation. With respect to VSAC half homography, on one hand the differences are almost undetectable when the images mostly overlap, as shown in Fig.2. On the other hand, MiHo reduces the distortions on the non-overlapping regions as the common scene area between the images is reduced, as shown in 3. By inspection, the differences between MiHo and VSAC half homograhy can mainly abducted to different camera center assumptions.
As detailed in [multiview], VSAC half homography corresponds to use conjugate rotations to generate synthetic views through fractional angles, implicitly assuming the same and general intrinsic camera matrix for both images. MiHo does not impose this constraint since in general , but indirectly includes a regularization factor towards affine transformations. Midpoints are preserved by similarity and affine transformation, being linear combinations, but not by general planar homographies.
MiHo homography strongly resembles affine or similarity transformations, imposing alternative constraints less restricted to the base homography. This model design is also reflected into different planes and inliers extracted by MOP and MOP+MiHo. To balance this additional restrain with respect to the original MOP, in MOP+MiHo the minimum requirement amount of inliers for increasing the failure counter defined by Eq. 4 has been relaxed after experimental validation from to .
@@ 3.2.4 Translation invariance
where is a fixed translation vector added to all the corresponding original midpoints . Overall, translating keypoints in the respective images have the effect of adding a fixed translation to the corresponding midpoint. Hence MiHo is invariant to translations since the normalization employed by the normalized 8-point algorithm is invariant to similarity transformations [multiview], including translations.